ABC corp acquired a machine on January 1, 2015 for $240,000. The machine has an estimated useful life of 4 years and a residual value of $40,000. ABC corp used the double-declining balance method of depreciation. What depreciation expense should ABC corp recognize for this machine in 2017?
Answer #1
Year Beginning Value Depreciation Expense Accumulated Depreciation Ending Value
2015 240,000 240,000 x 50% = 120,000 120,000 240,000-120,000 = 120,000
2016 120,000 120,000 x 50% = 60,000 120,000+60,000 = 180,000 240,000-180,000 = 60,000
2017 60,000 60,000 x 50%= 30,000 180,000+30,000 = 210,000 240,000-210,000 = 30,000

Double declining depreciation rate = 2 x 1/useful life

= 2 x 1/4

= 50%

Depreciation expense should ABC corp recognize for this machine in 2017 = $30,000

Kindly comment if you need further assistance. Thanks‼!